OpenStudy (anonymous):

can someone help me with this question 10y^2+9y-9 using grouping ac= and b= i have know clue what to do on this can someone help me with this please

OpenStudy (anonymous):

Are you looking to factor this?

OpenStudy (anonymous):

If so the factored form is: (2 y+3) (5 y-3)

OpenStudy (anonymous):

no just the ac= and that is what im stuck on to find the ac it that the number + the number or the number * the number

hero (hero):

ax^2 + bx + c

OpenStudy (anonymous):

ok so will a = 10y2 and c = -9

hero (hero):

The coefficients equal numbers only. Don't ever include the variables

hero (hero):

You must be confused again

OpenStudy (anonymous):

yes i am

hero (hero):

Find two numbers, m and n, that multiply to get ac, yet add to get b: mn = ac m + n = b

OpenStudy (anonymous):

ok so do i have to solve the problem like we did yesterday or do i just pull the 2 numbers out

hero (hero):

You always have to solve it like that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so will it be 9 and 1

hero (hero):

mn = -90 m + n = 9

hero (hero):

Funny thing is, we did something similar to this before using the exact same numbers.

hero (hero):

I see that you have forgotten them

hero (hero):

You only memorized your multiplication tables up to twelves, I bet.

OpenStudy (anonymous):

im sorry im bad at math not really i was cheating using a calculator

hero (hero):

I'm sorry to hear that.

OpenStudy (anonymous):

its all good

hero (hero):

15 - 6 = ?

OpenStudy (anonymous):

9

hero (hero):

Yep, so now you have what you need. Post your factorization here.

OpenStudy (anonymous):

10y^2+15y-9

hero (hero):

I thought you were getting the hang of it, but I see that you are not.

OpenStudy (anonymous):

ok so its 10y^2+15y-6

hero (hero):

You basically changed it to a different trinomial.

OpenStudy (anonymous):

o dang it well i thought i was getting it yesterday but i guess not

OpenStudy (anonymous):

and i have a test Monday on all of this stuff

hero (hero):

If you didn't write down the steps I showed you yesterday...

OpenStudy (anonymous):

i did wright them down

hero (hero):

10y^2+9y-9 10y^2 + 15y - 6y - 9 5y(2y + 3) -3(2y - 3) (2y + 3)(5y - 3)

hero (hero):

Notice that you only need to change the middle term, then you can start factoring: 9y = 15y - 6y

OpenStudy (anonymous):

o ok i forgot to put the 15 and the 6 together dang it

OpenStudy (anonymous):

i keep for getting that part

hero (hero):

Here's a page where you can practice. Scroll down. http://www.themathpage.com/alg/factoring-trinomials.htm

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

do you think you can help me more please

OpenStudy (anonymous):

12x^2-13x-4

hero (hero):

You're supposed to be practicing on your own.

hero (hero):

Can you at least come up with the equation for m and n?

OpenStudy (anonymous):

i can try to hold on

OpenStudy (anonymous):

ok so do have to change 12 because i cant find a number to make -13 and -4

hero (hero):

You're supposed to be solving this: mn = ac m + n = b In other words mn = (12)(-4) m + n = -13

hero (hero):

I usually multiply the a and the c and express it like this: mn = -48 m + n = -13

hero (hero):

You won't be able to get stuck like this on the test, so you better learn quick. Anyway, we have plenty to practice. Try doing the next step .

hero (hero):

This process of figuring out the m and the n is a mental process. You're supposed to ask yourself, "What are two numbers that multiply to get -48, yet add to get -13?".

hero (hero):

I just write it out explicitly so that you can see more easily what needs to be done.

OpenStudy (anonymous):

ok so its -16 and 3

OpenStudy (anonymous):

then the next step is 12x^2-16x-3x-4

hero (hero):

Did you use your calc to get that?

OpenStudy (anonymous):

i just divided 48 and 3and got it that way

hero (hero):

I see. Good job. Well...finish it. Wait. You have to make sure the middle two terms add to get -13. Double check your latest expression.

OpenStudy (anonymous):

12x^2-16x+3x-4

hero (hero):

There we go. If you figure out this next part, then you'll be well on your way.

hero (hero):

You have to factor the first two terms; then factor the last two terms. Then factor what's common to both.

OpenStudy (anonymous):

12x(x-16)+3(x-16) (x+16)(-3)

hero (hero):

Sigh

hero (hero):

12x^2-16x+3x-4 4x(3x - 4) + 1(3x - 4) (3x - 4)(4x + 1)

hero (hero):

Study that.

hero (hero):

Just a quick question....Do you know how to multiply binomials?

OpenStudy (anonymous):

not really

hero (hero):

That's where the problem lies.

hero (hero):

The intuition necessary to be able to factor trinomials comes from mastery of mulitplying binomials. You can't learn how to factor trinomials without first learning how to multiply binomials.

OpenStudy (anonymous):

my last math teacher didnt teach a lot he talked about other stuff like his family and the moon and stuff like that and now im having problems with this math class

hero (hero):

Well, let's learn how to multiply binomials real quick.

hero (hero):

Do you remember the distributive rule?

hero (hero):

Or rather, the distributive property?

hero (hero):

@jen89, are you still there?

OpenStudy (anonymous):

yes im still here sorry i had to do something really fast

OpenStudy (anonymous):

i dont realy remember that at all

hero (hero):

(a + c)b = ab + bc

hero (hero):

That's distributive rule.

hero (hero):

When multiplying binomials, you do a similar technique: (x + 3)(x + 4) = x(x + 4) + 3(x + 4)

hero (hero):

That's the first step

hero (hero):

x(x + 4) + 3(x + 4) = x^2 + 4x + 3x + 12

hero (hero):

x^2 + 4x + 3x + 12 = x^2 + 7x + 12

hero (hero):

after multiplying two binomials, you end up with a trinomial Thus (x + 3)(x + 4) = x^2 + 7x + 12

OpenStudy (anonymous):

ok i wrote the steps down

hero (hero):

Well, I was showing you step by step the process. Now, here's how I would usually introduce the steps to a student...in a more straight forward, but simpler form: (x + 3)(x + 4) = x(x + 4) + 3(x + 4) = x^2 + 4x + 3x + 12 = x^2 + 7x + 12

hero (hero):

Try to understand those steps.

hero (hero):

If you look closely, you'll notice that it's actually the reverse of factoring.

hero (hero):

Multiplying: (x + 3)(x + 5) = x(x + 5) + 3(x + 5) = x^2 + 5x + 3x + 15 = x^2 + 8x + 15 Factoring: x^2 + 8x + 15 = x^2 + 5x + 3x + 15 = x(x + 5) + 3(x + 5) = (x + 5)(x + 3)

hero (hero):

Multiplying binomials and factoring trinomials have an inverse relationship.

OpenStudy (anonymous):

ok i see that now when you put both together

hero (hero):

That is what you must master.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is there any way i can learn all of this before monday

hero (hero):

Yes, you can. all you need to do is study it. It is a template to use for other factoring methods. Basically, you just familiarize yourself with it. Try to think of it like getting to know someone you just met for the first time. When you meet someone for the first time and you're really interested in them, you try to learn as much as possible about them. You want to be a good friend to them and you may even want to have a mutual relationship with them. You should want to do something similar with learning this.

OpenStudy (anonymous):

ok thank you i will do my best learning all of this